Date:  Math 115  Name:
Start:  Derivative Gateway Test  ID #:
End:  Section:

Find the derivative of each of these. Calculators are not allowed. Passing is 6 out of 7. Allowed time: 15 minutes.

2. y = \sqrt{x^{2} − x + π}

24. y = (cos 2x)(ln \frac{1}{2}x)

41. y = \frac{e^{x}}{x^{2} − 3x − 1}

61. y = ^{7} \sqrt{(2x + 1)^{4}}

95. f(t) = 1 + \frac{e^{2t}}{2}

107. y = \sqrt{e^{x} + ln(2x)}

137. h(r) = ln(sin πr) BE SURE TO PICK UP YOUR STUDENT ID WHEN DONE!